OpenStudy (anonymous):

Which of the following is a solution of 9x2 - 5x = - 3?

OpenStudy (anonymous):

These are what the answer solutions have to look like x = -3 - i x = 3 + i x equals 3 plus i times the square root of 19 x equals negative 3 minus i times the square root of 19

OpenStudy (anonymous):

Add 3 to both sides, you get 9x^2 - 5x + 3 = 0 Now use the quadratic formula or completing the square to solve for x.

OpenStudy (thejax):

The formula,